OpenStudy (anonymous):

Trying to solve quadratic equation x2-2x-13=0

OpenStudy (anonymous):

you want the answer?

OpenStudy (mew55):

\[\frac{ +/- (b)\sqrt{b^2-4ac} }{ 2a }\] ur values r b=-2 a=1 c=-13 keep in mind, on the first b value with the plus and minus, the -2 becomes a positive 2. inside the square root will look like 4-4(1)(-13) and the denominator will look like 2(1).

OpenStudy (mew55):

the rest is up to u to solve

OpenStudy (anonymous):

alright. I'll give you the answer.

OpenStudy (anonymous):

1 +/- radical 14

OpenStudy (mew55):

yep that is right. i was just about to solve it XD. the set up i inputed is right?

OpenStudy (anonymous):

yes just this part: